How to solve
1.
2x+3y=13
3x+5y=40
2. 3x+5y=7
4x+6y=12

Answer by Cromlix(4381)

Hi there,
2x+3y=13........Eq(1)
3x+5y=40........Eq(2)
Multiply Eq(1) by 3
Multiply Eq(2) by 2
6x + 9y = 39.....Eq(1)
6x +10y = 80.....Eq(2)
Subtract Eq(1) from Eq(20
......y = 41
Substitute y = 41 into Eq(10
2x + 3y = 13
2x + 123 = 13
2x = 13 - 123
2x = -110
x = -55 [x = -55 and y = 41]
..............
3x+5y=7.........Eq(1)
4x+6y=12........Eq(2)
Multiply Eq(1) by 4
Multiply Eq(2) by 3
12x + 20y = 28......Eq(1)
12x + 18y = 36......Eq(2)
Subtract Eq(2) from Eq(1)
.......2y = -8
........y = -4
Substitute y = -4 into Eq(2)
4x + 6y = 12.......Eq(2)
4x + (-24) = 12
4x - 24 = 12
4x = 12 + 24
4x = 36
x = 9 [x = 9 and y = -4]
Hope this helps :-)